# Degrees of interior polynomials and parking function enumerators

### v2

@@ Abstract.
The interior polynomial of a directed graph is defined as the $h^{*}$-polynomial of the graph’s (extended) root polytope, and it displays several attractive properties. Here we express its degree in terms of the minimum cardinality of a directed join, and give a formula for the leading coefficient. We present natural generalizations of these results to oriented regular matroids; in the process we also give a facet description for the extended root polytope of a regular oriented matroid.
@@ 1.1. Interior polynomials
The first type of polynomial we deal with is the interior polynomial of a directed graph. This notion generalizes some well-studied graph invariants, most notably the specialization $T(x,1)$ of the Tutte polynomial, as well as its extension to hypergraphs [10]. It is defined as the $h^{*}$-vector of the (extended) root polytope
@@
associated to the digraph $G=(V,E)$. Please see Section 2 for detailed definitions. In this paper we show that the degree and leading coefficient of the interior polynomial have meaningful connections to the graph structure.
@@
Recall that by a theorem of Lucchesi and Younger [16], the quantity $\nu(G)$ above is also the maximal number of edge-disjoint directed cuts in $G$. Furthermore, if the underlying undirected graph of $G$ is 2-edge connected, then $\nu(G)$ is the minimal number of edges whose reversal yields a strongly connected orientation of $G$ [8, Proposition 9.7.1].
@@
The leading coefficient of $I_{G}$ equals the number of vectors that can be obtained as the net degree vector of a minimal cardinality dijoin in $G$.
@@
Figure 1 shows a directed graph $G$ whose interior polynomial is $I_{G}(x)=4x^{2}+3x+1$. This can be checked directly in various ways, one of which is mentioned in the next Remark. Note that the coefficient of $x$ indeed agrees with $|E|-|V|+1=3$. As to the leading coefficient, first notice that the edge set of $G$ in this case is the disjoint union of five directed cuts, whence $\nu(G)\geq 5$. In fact $\nu(G)=5$ since it is easy to find dijoins of cardinality $5$ (actually, there are $18$ of them). Thus Theorem 1.1 confirms that $\deg I_{G}=8-1-5=2$ and by Theorem 1.3, the leading coefficient should indeed be $4$ because the minimal dijoins induce the following four net degree vectors:
@@
Let us explain a simple case that inspired Theorem 1.1. The interior polynomial was first introduced by Kálmán [10], for hypergraphs. By the main result of [11], that definition corresponds to the situation when $G=(U,W;E)$ is a bipartite graph with classes $U$ and $W$, and each edge is oriented toward $W$. (The graph $G$ of Example 1.4 is of this type with $|U|=|W|=4$, $\nu(G)=5$, and $\deg I_{G}=2$. In particular, its interior polynomial may be evaluated by using [10, Definition 5.3].)
The hypergraphical approach to the interior polynomial $I_{G}$ easily yields (see [10, Proposition 6.1]) that (for graphs $G$ as above)
@@
By analogy to directed graphs, one can introduce the (extended) root polytope of a regular oriented matroid and define the interior polynomial as its $h^{*}$-vector. (For details, see Section 4.) It turns out that all arguments readily generalize to this case, and one is able to obtain analogous statements regarding the degree and leading coefficient.
@@
Let $M$ be a regular oriented matroid of rank $r$. Then the degree of the interior polynomial $I_{M}$ of $M$ is $r-\nu(M)$, where $\nu(M)=\min\{|K|\mid K\text{ is a dijoin of }M\}$.
@@
Take an arbitrary totally unimodular representing matrix $A=[\mathbf{a}_{i}]_{i\in E}$ for the regular oriented matroid $M$. The leading coefficient of $I_{M}$ equals the number of vectors that can be obtained as $\sum_{i\in K}\mathbf{a}_{i}$, where $K$ is a minimal cardinality dijoin of $M$.
Even though Theorems 1.6 and 1.7 imply Theorems 1.1 and 1.3, respectively, we include separate proofs for the benefit of those readers who are less interested in oriented matroids. A facet description of the extended root polytope is provided in Proposition 4.1. We remark that D’Ali, Juhnke-Kubitzke and Koch [7] found a similar facet description in a special case, namely the one that corresponds to the symmetric edge polytope of a graph.
It is interesting to compare interior polynomials for different orientations of a given undirected graph. In an earlier paper [12] we highlighted a special case of this problem. Namely, we conjectured that for a bipartite graph, among its so-called semi-balanced orientations (see Section 2), the one in which all edges point toward the same class has a coefficientwise minimal interior polynomial. (There are of course two such orientations but their extended root polytopes are isometric and thus their interior polynomials identical.) Semi-balanced orientations are relevant because their root polytopes are the facets of the symmetric edge polytope of the bipartite graph. Here we prove a weakened version of this conjecture, addressing only the degree: We show that, in fact among all orientations of a bipartite graph, the one with each edge pointing toward one of the classes has an interior polynomial of minimal degree. The analogous problem remains open for non-bipartite graphs and for regular matroids, where we do not even have a candidate for the degree-minimizing orientation. We do however conjecture that in these broader cases, too, there is an orientation that coefficientwise minimizes the interior polynomial.
1.2. Greedoid polynomials
The other main object of this paper is the greedoid polynomial, introduced by Björner, Korte, and Lovász [5]. It is set up in such a way that interesting information is contained in its lowest-degree term. (The degree of the leading term depends only weakly on the greedoid, namely it is the size of the ground set minus the rank. For exact definitions, see Section 5.)
@@ 1.2. Greedoid polynomials
The degree of the parking function enumerator of a connected Eulerian digraph $G$ (with any root) is equal to $|E(G)|-|V(G)|+1-{\rm minfas}(G)$, where ${\rm minfas}(G)$ denotes the minimal cardinality of a feedback arc set in $G$.
@@
By [5, Theorem 6.10], the constant term of the greedoid polynomial of a (root-connected) rooted digraph $G$ is zero if and only if $G$ contains a directed cycle. Theorem 1.10 strengthens this statement. Indeed, (for an $s$-root-connected digraph $G$) we have ${\rm minfas}(G,s)>0$ if and only if $G$ contains a directed cycle.
@@
The paper is structured as follows. Section 2 contains our results on the interior polynomial in the most ‘classical’ case of directed graphs, including the proofs of Theorems 1.1 and 1.3. Section 3 discusses the comparison of interior polynomials for different orientations of the same graph. In Section 4 we generalize Section 2 to regular oriented matroids and prove Theorems 1.6 and 1.7. Finally, in Section 5 we turn to parking function enumerators and greedoid polynomials, and prove Theorems 1.8, 1.12, and 1.10.
@@ 1.4. Graph notation
Let $G=(V,E)$ be a directed graph (abbreviated to digraph throughout). A non-empty set of edges $C^{*}\subseteq E(G)$ is called a cut if there is a partition $V_{0}\sqcup V_{1}=V(G)$ of the vertex set such that $C^{*}$ is the set of edges connecting a vertex of $V_{0}$ and a vertex of $V_{1}$. The sets $V_{0}$ and $V_{1}$ are called the shores of the cut. Note that in a connected graph, a cut uniquely determines its two shores. A cut $C^{*}$ is called elementary if it is minimal among cuts with respect to containment; this is equivalent to the condition that $V_{0}$ and $V_{1}$ both span (weakly) connected subgraphs. A cut $C^{*}$ is directed if either each edge of $C^{*}$ leads from $V_{0}$ to $V_{1}$, or each edge of $C^{*}$ leads from $V_{1}$ to $V_{0}$.
@@
A set of edges $K\subseteq E(G)$ is called a directed join or dijoin if for any directed cut $C^{*}$ of $G$ we have $C^{*}\cap K\neq\emptyset$. We denote the cardinality of a minimal dijoin of $G$ by $\nu(G)$. A set of edges $F\subseteq E(G)$ is called a feedback arc set if for any directed cycle $C$ we have $C\cap F\neq\emptyset$. We denote the minimal cardinality of a feedback arc set of $G$ by ${\rm minfas}(G)$.
Let $G$ be a digraph with a fixed root vertex $s$. The graph $G$ is said to be $s$-root-connected if each vertex is reachable along a directed path from $s$. We call a subgraph $A$ an arborescence rooted at $s$ if its underlying undirected graph is a tree and it is $s$-root-connected. A spanning arborescence is an arborescence that contains each vertex of $G$. We will denote the set of spanning arborescences of $G$, rooted at $s$, by ${\rm Arb}(G,s)$. We note that a digraph $G$, rooted at $s$, has a spanning arborescence (rooted at $s$) if and only if it is $s$-root-connected.
For a digraph $G=(V,E)$, and an edge set $S\subseteq E$, we denote by $G[S]$ the digraph with vertex set $V$ and edge set $S$.
@@ 2. The degree of the interior polynomial of a digraph
Let $G=(V,E)$ be a directed graph. To an edge $e=\overrightarrow{th}\in E$, let us associate the vector $\mathbf{x}_{e}=\mathbf{1}_{h}-\mathbf{1}_{t}\in\mathbb{R}^{V}$. (Here $t,h\in V$ and $\mathbf{1}_{t},\mathbf{1}_{h}\in\mathbb{R}^{V}$ are the corresponding generators.)
@@
Let us recall the well known facts that for $G=T$ a tree (or forest), both $\mathcal{Q}_{T}$ and $\tilde{\mathcal{Q}}_{T}$ are simplices, moreover, they are unimodular with respect to the lattice $\mathbb{Z}^{V}$. See, for example, [13] for details and proofs.
@@
Before proving Theorem 1.1, we examine root polytopes in more detail. If $G$ is connected, then the dimension of $\mathcal{Q}_{G}$ is either $|V|-1$ or $|V|-2$. It is $|V|-2$ if and only if $G$ satisfies the following condition [13].
@@
We note that an alternative characterization of semi-balanced digraphs is that each cycle has the same number of edges going in the two directions around the cycle. This description was given as the definition in [13].
Now let us turn to the extended root polytope $\tilde{\mathcal{Q}}_{G}$. Note that if $G$ has a directed cycle $C$ then $\frac{1}{|C|}\sum_{e\in C}\mathbf{x}_{e}=\mathbf{0}$, that is, $\mathbf{0}$ is a point of $\mathcal{Q}_{G}$. Hence if $G$ is not acyclic, then $\mathcal{Q}_{G}=\tilde{\mathcal{Q}}_{G}$.
@@
Hence if we defined the interior polynomial using $\mathcal{Q}_{G}$ instead of the extended root polytope, that would only make a difference for acyclic, but not semi-balanced graphs. However, Theorem 1.1 would not hold if the $h^{*}$-polynomial of $\mathcal{Q}_{G}$ replaced that of ${\tilde{\mathcal{Q}}}_{G}$.
@@
As an alternative way to think of the extended root polytope, note that loop edges $f$ have $\mathbf{x}_{f}=\mathbf{0}$. Hence ${\tilde{\mathcal{Q}}}_{G}=\mathcal{Q}_{G\cup f}$, where $f$ is a loop edge attached to any vertex of $G$.
@@
We remark that any admissible layering of a semi-balanced graph is actually a layering in the sense of Definition 2.4, so we have not introduced anything new into that context.
@@
$\tilde{\mathcal{Q}}_{G}=\left\{\mathbf{x}\in\mathbb{R}^{V}\middle|\begin{array}[]{cl}f_{C^{*}}(\mathbf{x})\geq 0&\text{for all elementary directed cuts }C^{*}\text{ of }G\\ \ell\cdot\mathbf{x}\leq 1&\text{for all admissible layerings $\ell$ of $G$}\\ \mathbf{1}\cdot\mathbf{x}=0&\end{array}\right\}.$
Here $\mathbf{1}=\sum_{v\in V}\mathbf{1}_{v}\in\mathbb{R}^{V}$. Note that if $\ell$ is an admissible layering then so is $\ell+m\cdot\mathbf{1}$ for all $m\in\mathbb{Z}$. As is obvious from the Proposition, members of such equivalence classes of admissible layerings describe the same constraint for $\tilde{\mathcal{Q}}_{G}$. If we chose a representative of each class, for example by assuming that $\ell(v_{0})=0$ for some fixed vertex $v_{0}$, then the number of constraints would become formally finite. (Indeed, there are only finitely many ways to choose the connected subgraph whose edges $\overrightarrow{th}$ satisfy $\ell(h)-\ell(t)=1$, and with $\ell(v_{0})=0$ that subgraph already determines $\ell$.)
Proposition 2.8 contains some previously published special cases, such as [11, Proposition 3.6] and [9, Theorem 3.1]. A very similar facet description of $\mathcal{Q}_{G}$ was recently obtained by Numata et al. [17].
@@
First note that each vector $\mathbf{x}\in{\tilde{\mathcal{Q}}}_{G}$ satisfies the conditions of the right hand side. Clearly $\mathbf{1}\cdot\mathbf{x}_{e}=0$ for each $e\in E(G)$ and $\mathbf{1}\cdot\mathbf{0}=0$, whence $\mathbf{1}\cdot\mathbf{x}=0$ for each $\mathbf{x}\in\tilde{\mathcal{Q}}_{G}$. Similarly for any admissible layering $\ell$, by definition we have $\ell\cdot\mathbf{x}_{e}\leq 1$ for each $e\in E$, which in addition to the obvious $\ell\cdot\mathbf{0}=0$ implies $\ell\cdot\mathbf{x}\leq 1$ for each $\mathbf{x}\in{\tilde{\mathcal{Q}}}_{G}$. Finally let $C^{*}$ be any directed cut. Then by definition, the induced functional $f_{C^{*}}$ is such that $f_{C^{*}}(\mathbf{x}_{e})=0$ for $e\notin C^{*}$ and $f_{C^{*}}(\mathbf{x}_{e})=1$ for $e\in C^{*}$; moreover $f_{C^{*}}(\mathbf{0})=0$. Hence each $\mathbf{x}\in{\tilde{\mathcal{Q}}}_{G}$ satisfies $f_{C^{*}}(\mathbf{x})\geq 0$.
Conversely, we show that any element $\mathbf{x}$ of the right hand side belongs to ${\tilde{\mathcal{Q}}}_{G}$. From the first part of the proof, and the fact that ${\tilde{\mathcal{Q}}}_{T}$ is a $(|V|-1)$-dimensional simplex for any spanning tree $T\subseteq G$, it is already clear that $\dim{\tilde{\mathcal{Q}}}_{G}=|V|-1$. The condition $\mathbf{1}\cdot\mathbf{x}=0$ says that $\mathbf{x}$ is in the same hyperplane of $\mathbb{R}^{V}$ as ${\tilde{\mathcal{Q}}}_{G}$; in the rest of the proof we work relative to this hyperplane.
@@
The facet $F$ needs to contain $|V|-1$ affine independent vertices of ${\tilde{\mathcal{Q}}}_{G}$. If $F$ does not contain $\mathbf{0}$, then those are $|V|-1$ affine independent vertices of type $\mathbf{x}_{e}$, which means that the corresponding edges $e$ form a spanning tree $T$. The tree $T$, along with the orientations of its edges, determines a layering $\ell\colon V\to\mathbb{Z}$ so that $\ell(\mathbf{x}_{e})=1$ for all $e\in T$. This layering is necessarily admissible, for otherwise there would be an edge $f$ with $\ell(\mathbf{x}_{f})\geq 2$, meaning that the vertices $\mathbf{0}$ and $\mathbf{x}_{f}$ lie on opposite sides of the hyperplane of $F$. Therefore the hyperplane of the facet is $\{\mathbf{x}\mid\ell\cdot\mathbf{x}=1\}$ for the admissible layering $\ell$, and as $\ell\cdot\mathbf{0}=0$, the polytope $\tilde{\mathcal{Q}}_{G}$ lies in the half-space $\{\mathbf{p}\mid\ell\cdot\mathbf{p}\leq 1\}$, whence indeed the hyperplane of $F$ does not separate $\mathbf{x}$ from ${\tilde{\mathcal{Q}}}_{G}$.

### v3

@@ Abstract.
The interior polynomial of a directed graph is defined as the $h^{*}$-polynomial of the graph’s (extended) root polytope, and it displays several attractive properties. Here we express its degree in terms of the minimum cardinality of a directed join, and give a formula for the leading coefficient. We present natural generalizations of these results to oriented regular matroids; in the process we also give a facet description for the extended root polytope of an oriented regular matroid.
@@ 1.1. Interior polynomials
The first type of polynomial we deal with is the interior polynomial of a directed graph. This notion generalizes some well-studied graph invariants, most notably the specialization $T(x,1)$ of the Tutte polynomial, as well as its extension to hypergraphs [12]. It is defined as the $h^{*}$-vector of the (extended) root polytope [10]
@@
associated to the digraph $G=(V,E)$. We denote the interior polynomial of $G$ by $I_{G}$. When $G$ is a bidirected graph, ${\tilde{\mathcal{Q}}}_{G}$ is known as the symmetric edge polytope of the underlying undirected graph. Please see Section 2 for detailed definitions. In this paper we show that the degree and leading coefficient of the interior polynomial have meaningful connections to the graph structure.
@@
Recall that by a theorem of Lucchesi and Younger [18], the quantity $\nu(G)$ above is also the maximal number of edge-disjoint directed cuts in $G$. Furthermore, if the underlying undirected graph of $G$ is 2-edge connected, then $\nu(G)$ is the minimum number of edges whose reversal yields a strongly connected orientation of $G$ [9, Proposition 9.7.1].
@@
The leading coefficient of $I_{G}$ equals the number of vectors that can be obtained as the net degree vector of a minimum cardinality dijoin in $G$.
@@
Remark 1.4.
In [10], Higashitani proved that the (extended) root polytope of a strongly connected digraph is reflexive, whence its interior polynomial has degree $|V|-1$ and it is palindromic. Thus for a strongly connected digraph, the leading coefficient is also $1$. We can recover these results from Theorems 1.1 and 1.3. Indeed, for a strongly connected digraph $G$ we have $\nu(G)=0$, and the unique net degree vector of minimal cardinality dijoins is the zero vector.
@@
Figure 1 shows a directed graph $G$ whose interior polynomial is $I_{G}(x)=4x^{2}+3x+1$. This can be checked directly in various ways, one of which is mentioned in the next Remark. Note that the coefficient of $x$ indeed agrees with $|E|-|V|+1=3$. As to the leading coefficient, first notice that the edge set of $G$ is the disjoint union of five directed cuts, whence $\nu(G)\geq 5$. In fact $\nu(G)=5$ since it is easy to find dijoins of cardinality $5$ (actually, there are $18$ of them). Thus Theorem 1.1 confirms that $\deg I_{G}=8-1-5=2$ and by Theorem 1.3, the leading coefficient should indeed be $4$ because the minimum cardinality dijoins induce the following four net degree vectors:
@@
Let us explain a simple case that inspired Theorem 1.1. The interior polynomial was first introduced by Kálmán [12], for hypergraphs. By the main result of [13], that definition corresponds to the situation when $G=(U,W;E)$ is a bipartite graph with classes $U$ and $W$, and each edge is oriented toward $W$. (The graph $G$ of Example 1.5 is of this type with $|U|=|W|=4$, $\nu(G)=5$, and $\deg I_{G}=2$. In particular, its interior polynomial may be evaluated by using [12, Definition 5.3].)
The hypergraphical approach to the interior polynomial $I_{G}$ easily yields (see [12, Proposition 6.1]) that (for graphs $G$ as above)
@@
By analogy to directed graphs, one can introduce the (extended) root polytope of an oriented regular matroid and define the interior polynomial as its $h^{*}$-vector. (For details, see Section 4.) It turns out that all arguments readily generalize to this case, and one is able to obtain analogous statements regarding the degree and leading coefficient.
@@
Let $M$ be an oriented regular matroid of rank $r$. Then the degree of the interior polynomial $I_{M}$ of $M$ is $r-\nu(M)$, where $\nu(M)=\min\{|K|\mid K\text{ is a dijoin of }M\}$.
@@
Take an arbitrary totally unimodular representing matrix $A=[\mathbf{a}_{i}]_{i\in E}$ for the oriented regular matroid $M$. The leading coefficient of $I_{M}$ equals the number of vectors that can be obtained as $\sum_{i\in K}\mathbf{a}_{i}$, where $K$ is a minimum cardinality dijoin of $M$.
Even though Theorems 1.7 and 1.8 imply Theorems 1.1 and 1.3, respectively, we include separate proofs for the benefit of those readers who are less interested in oriented matroids. A facet description of the extended root polytope is provided in Proposition 4.1. We remark that D’Alì, Juhnke-Kubitzke and Koch [8] found a similar facet description in a special case, namely the one that corresponds to the symmetric edge polytope of a graph.
It is interesting to compare interior polynomials for different orientations of a given undirected graph. In an earlier paper [14] we highlighted a special case of this problem. Namely, we conjectured that for a bipartite graph, among its so-called semi-balanced orientations (see Section 2), the one in which all edges point toward the same class has a coefficientwise minimal interior polynomial. (There are of course two such orientations but their extended root polytopes are isometric and thus their interior polynomials identical.) Semi-balanced orientations are relevant because their root polytopes are the facets of the symmetric edge polytope of the bipartite graph. Here we prove a weakened version of this conjecture, addressing only the degree: We show that, in fact among all orientations of a bipartite graph, the one with each edge pointing toward one of the classes has an interior polynomial of minimum degree. It remains an open problem which orientations minimize the degree of the interior polynomial among non-bipartite graphs. On the other hand, in section 3 we show an example that among non-bipartite graphs, there does not always exist an orientation coefficientwise minimizing the interior polynomial.
The analogous questions regarding maxima are settled as follows: the degree of the interior polynomial is maximized exactly by the strongly connected orientations (at least as long as the graph is connected and without bridge edges, see section 3); coefficientwise maximizing orientations do not always exist.
1.2. Greedoid polynomials and parking function enumerators
The other main object of this paper is the greedoid polynomial, introduced by Björner, Korte, and Lovász [6]. A special case of this notion is equivalent, via a change of variable, to parking function enumerators. Greedoid polynomials are set up in such a way that interesting information is contained in their lowest-degree term. (The degree of the leading term depends only weakly on the greedoid, namely it is the size of the ground set minus the rank. For exact definitions, see Section 5.)
@@ 1.2. Greedoid polynomials and parking function enumerators
The degree of the parking function enumerator of a connected Eulerian digraph $G$ (with any root) is equal to $|E(G)|-|V(G)|+1-{\rm minfas}(G)$, where ${\rm minfas}(G)$ denotes the minimum cardinality of a feedback arc set in $G$.
@@
By [6, Theorem 6.10], the constant term of the greedoid polynomial of a (root-connected) rooted digraph $G$ is zero if and only if $G$ contains a directed cycle. Theorem 1.11 strengthens this statement. Indeed, (for an $s$-root-connected digraph $G$) we have ${\rm minfas}(G,s)>0$ if and only if $G$ contains a directed cycle.
@@
The paper is structured as follows. Section 2 contains our results on the interior polynomial in the most ‘classical’ case of directed graphs, including the proofs of Theorems 1.1 and 1.3. Section 3 discusses the comparison of interior polynomials for different orientations of the same graph. In Section 4 we generalize Section 2 to oriented regular matroids and prove Theorems 1.7 and 1.8. Finally, in Section 5 we turn to parking function enumerators and greedoid polynomials, and prove Theorems 1.9, 1.13, and 1.11.
@@ 1.4. Graph notation
Let $G=(V,E)$ be a directed graph (abbreviated to digraph throughout). We allow loops and multiple edges. For an edge set $S\subseteq E$, we denote by $G[S]$ the digraph with vertex set $V$ and edge set $S$.
By a subgraph, in this paper we will mean an object as above. That is, all vertices are automatically included, even if they become isolated in the subgraph. We call a subgraph $G[S]$ spanning if it has the same number of connected components as $G$. (We note that the term spanning subgraph is sometimes used in a different sense. Our usage correlates with spanning sets in matroid theory.) In particular for a subgraph of a connected graph, being connected and spanning mean the same thing. We will mostly use the latter term for trees and the former otherwise. When we want to emphasize that connectedness is meant without regard to edge orientations, we will speak of weakly connected (sub)graphs.
A non-empty set of edges $C^{*}\subseteq E(G)$ is called a cut if there is a partition $V_{0}\sqcup V_{1}=V(G)$ of the vertex set such that $C^{*}$ is the set of edges connecting a vertex of $V_{0}$ and a vertex of $V_{1}$. The sets $V_{0}$ and $V_{1}$ are called the shores of the cut. Note that in a connected graph, a cut uniquely determines its two shores. A cut $C^{*}$ is called elementary if it is minimal among cuts with respect to containment; this is equivalent to the condition that $E(G)-C^{*}$, as a subgraph, has exactly two (weakly) connected components. A cut $C^{*}$ is directed if either each edge of $C^{*}$ leads from $V_{0}$ to $V_{1}$, or each edge of $C^{*}$ leads from $V_{1}$ to $V_{0}$.
@@
A set of edges $K\subseteq E(G)$ is called a directed join or dijoin if for any directed cut $C^{*}$ of $G$ we have $C^{*}\cap K\neq\emptyset$. We denote the minimum cardinality of a dijoin of $G$ by $\nu(G)$. A set of edges $F\subseteq E(G)$ is called a feedback arc set if for any directed cycle $C$ we have $C\cap F\neq\emptyset$. We denote the minimum cardinality of a feedback arc set of $G$ by ${\rm minfas}(G)$.
Let $G$ be a digraph with a fixed root vertex $s$. The graph $G$ is said to be $s$-root-connected if each vertex is reachable along a directed path from $s$. We call a subgraph an arborescence rooted at $s$ if it consists of an $s$-root-connected tree and some isolated points. A spanning arborescence is an arborescence without isolated points. We will denote the set of spanning arborescences of $G$, rooted at $s$, by ${\rm Arb}(G,s)$. We note that a digraph $G$, rooted at $s$, has a spanning arborescence (rooted at $s$) if and only if it is $s$-root-connected.
@@ 2. The degree of the interior polynomial of a digraph
Let $G=(V,E)$ be a directed graph. To an edge $e=\overrightarrow{th}\in E$, let us associate the vector $\mathbf{x}_{e}=\mathbf{1}_{h}-\mathbf{1}_{t}\in\mathbb{R}^{V}$. (Here $t,h\in V$ and $\mathbf{1}_{t},\mathbf{1}_{h}\in\mathbb{R}^{V}$ are the corresponding generators.) The first of the following two notions was introduced by Higashitani while studying smooth Fano polytopes [10].
@@
Let us recall the following well known facts. See, for example, [15, Lemma 3.5 and Corollary 3.6] for details and proofs.
Lemma 2.2.
For $G=T$ a tree (or forest), both $\mathcal{Q}_{T}$ and $\tilde{\mathcal{Q}}_{T}$ are simplices, moreover, they are unimodular with respect to the lattice $\mathbb{Z}^{V}$.
@@
Before proving Theorem 1.1, we examine root polytopes in more detail. If $G$ is connected, then by [10, Proposition 1.3] and [15, Corollary 3.2] the dimension of $\mathcal{Q}_{G}$ is either $|V|-1$ or $|V|-2$. It is $|V|-2$ if and only if $G$ satisfies the following condition.
@@
We note that an alternative characterization of semi-balanced digraphs [15, Theorem 2.6] is that each cycle has the same number of edges going in the two directions around the cycle. This description was given as the definition in [15].
Now let us turn to the extended root polytope $\tilde{\mathcal{Q}}_{G}$. As an alternative way to think of it, note that loop edges $f$ have $\mathbf{x}_{f}=\mathbf{0}$. Hence ${\tilde{\mathcal{Q}}}_{G}=\mathcal{Q}_{G\cup f}$, where $f$ is a loop edge attached to any vertex of $G$. As a digraph with a loop is never semibalanced, this implies the following.
Proposition 2.6.
For a connected graph $G$, the dimension of ${\tilde{\mathcal{Q}}}_{G}$ is $|V|-1$.
Note that if $G$ has a directed cycle $C$ then $\frac{1}{|C|}\sum_{e\in C}\mathbf{x}_{e}=\mathbf{0}$, that is, $\mathbf{0}$ is a point of $\mathcal{Q}_{G}$. Hence if $G$ is not acyclic, then $\mathcal{Q}_{G}=\tilde{\mathcal{Q}}_{G}$.
@@
Hence if we defined the interior polynomial using $\mathcal{Q}_{G}$ instead of the extended root polytope, that would only make a difference for acyclic, but not semi-balanced graphs. However, in that class we do find examples which show that Theorem 1.1 would not hold if the $h^{*}$-polynomial of $\mathcal{Q}_{G}$ replaced that of ${\tilde{\mathcal{Q}}}_{G}$:
@@
It is not hard to show that any admissible layering of a semi-balanced graph is actually a layering in the sense of Definition 2.5, so we have not introduced anything new into that context. We will not rely on this fact later.
@@
(2.2) $\tilde{\mathcal{Q}}_{G}=\left\{\mathbf{x}\in\mathbb{R}^{V}\middle|\begin{array}[]{cl}f_{C^{*}}(\mathbf{x})\geq 0&\text{for all elementary directed cuts }C^{*}\text{ of }G\\ \ell\cdot\mathbf{x}\leq 1&\text{for all admissible layerings $\ell$ of $G$}\\ \mathbf{1}\cdot\mathbf{x}=0&\end{array}\right\}.$
Here $\mathbf{1}=\sum_{v\in V}\mathbf{1}_{v}\in\mathbb{R}^{V}$. It will be clear from the proof below that $f_{C^{*}}(\mathbf{x})\geq 0$, whenever $\mathbf{x}\in\tilde{\mathcal{Q}}_{G}$, holds for non-elementary directed cuts $C^{*}$, too.
Note that if $\ell$ is an admissible layering then so is $\ell+m\cdot\mathbf{1}$ for all $m\in\mathbb{Z}$. As is obvious from the proposition, members of such equivalence classes of admissible layerings describe the same constraint for $\tilde{\mathcal{Q}}_{G}$. If we chose a representative of each class, for example by assuming that $\ell(v_{0})=0$ for some fixed vertex $v_{0}$, then the number of constraints would become formally finite. (Indeed, there are only finitely many ways to choose the connected subgraph whose edges $\overrightarrow{th}$ satisfy $\ell(h)-\ell(t)=1$, and with $\ell(v_{0})=0$ that subgraph already determines $\ell$.)
Corollary 2.11.
For any connected digraph $G$, the facets of $\tilde{\mathcal{Q}}_{G}$ are as follows:
  1. (1)
    $C^{*}\mapsto\{\mathbf{x}\in\tilde{\mathcal{Q}}_{G}\mid f_{C^{*}}(\mathbf{x})=0\}$ gives a one-to-one correspondence between the elementary directed cuts of $G$ and facets containing $\mathbf{0}$.
  2. (2)
    $\ell\mapsto\{\mathbf{x}\in\tilde{\mathcal{Q}}_{G}\mid\ell\cdot\mathbf{x}=1\}$ induces a bijection between the equivalence classes of admissible layerings described above and facets not containing $\mathbf{0}$.
Proposition 2.10 contains some previously published special cases, such as [13, Proposition 3.6] and [11, Theorem 3.1]. A facet description of $\mathcal{Q}_{G}$, very similar to Corollary 2.11, was recently obtained by Numata et al. [19].
@@
First note that each vector $\mathbf{x}\in{\tilde{\mathcal{Q}}}_{G}$ satisfies the conditions of the right hand side of (2.2). Clearly $\mathbf{1}\cdot\mathbf{x}_{e}=0$ for each $e\in E(G)$ and $\mathbf{1}\cdot\mathbf{0}=0$, whence $\mathbf{1}\cdot\mathbf{x}=0$ for each $\mathbf{x}\in\tilde{\mathcal{Q}}_{G}$. Similarly for any admissible layering $\ell$, by definition we have $\ell\cdot\mathbf{x}_{e}\leq 1$ for each $e\in E$, which in addition to the obvious $\ell\cdot\mathbf{0}=0$ implies $\ell\cdot\mathbf{x}\leq 1$ for each $\mathbf{x}\in{\tilde{\mathcal{Q}}}_{G}$. Finally let $C^{*}$ be any directed cut. Then by definition, the induced functional $f_{C^{*}}$ is such that $f_{C^{*}}(\mathbf{x}_{e})=0$ for $e\notin C^{*}$ and $f_{C^{*}}(\mathbf{x}_{e})=1$ for $e\in C^{*}$; moreover $f_{C^{*}}(\mathbf{0})=0$. Hence each $\mathbf{x}\in{\tilde{\mathcal{Q}}}_{G}$ satisfies $f_{C^{*}}(\mathbf{x})\geq 0$.
Conversely, we show that any element $\mathbf{x}$ of the right hand side belongs to ${\tilde{\mathcal{Q}}}_{G}$. Proposition 2.6 claimed that $\dim{\tilde{\mathcal{Q}}}_{G}=|V|-1$. The condition $\mathbf{1}\cdot\mathbf{x}=0$ says that $\mathbf{x}$ is in the same hyperplane of $\mathbb{R}^{V}$ as ${\tilde{\mathcal{Q}}}_{G}$; in the rest of the proof we work relative to this hyperplane.
@@
The facet $F$ needs to contain $|V|-1$ affine independent vertices of ${\tilde{\mathcal{Q}}}_{G}$. If $F$ does not contain $\mathbf{0}$, then those are $|V|-1$ affine independent vertices of type $\mathbf{x}_{e}$, which means that the corresponding edges $e$ form a spanning tree $T$. The tree $T$, along with the orientations of its edges, determines a layering $\ell\colon V\to\mathbb{Z}$ so that $\ell\cdot\mathbf{x}_{e}=1$ for all $e\in T$. This layering is necessarily admissible, for otherwise there would be an edge $f$ with $\ell\cdot\mathbf{x}_{f}\geq 2$, meaning that the vertices $\mathbf{0}$ and $\mathbf{x}_{f}$ lie on opposite sides of the hyperplane of $F$. Therefore the hyperplane of the facet is $\{\mathbf{x}\mid\ell\cdot\mathbf{x}=1\}$ for the admissible layering $\ell$, and as $\ell\cdot\mathbf{0}=0$, the polytope $\tilde{\mathcal{Q}}_{G}$ lies in the half-space $\{\mathbf{p}\mid\ell\cdot\mathbf{p}\leq 1\}$, whence indeed the hyperplane of $F$ does not separate $\mathbf{x}$ from ${\tilde{\mathcal{Q}}}_{G}$.